OpenStudy (anonymous):

simplest slope for the equation -(1/2)x-y=(3/4)

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

remember the equation for a line is: y=mx+b y=1/2x +3/4 so m=1/2

OpenStudy (anonymous):

is that what you want?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok